a man was 4 times less than 1 year as old as his son 5years ago.in 4 years time,the man wil b 2time plus4 years as old the son.find their present age
Answer by ankor@dixie-net.com(22740)

let m = man's present age
let s = son's present age
:
Write an equation for each statement; simplify as much as possible
:
a man was 4 times less than 1 year as old as his son 5years ago.
m - 5 = 4(s-5) - 1
m - 5 = 4s - 20 - 1
m = 4s - 20 - 1 + 5
m = 4s - 16
:
in 4 years time, the man wil b 2time plus4 years as old the son.
m + 4 = 2(s+4) + 4
m + 4 = 2s + 8 + 4
m + 4 = 2s + 12
m = 2s + 12 - 4
m = 2s + 8
:
m = m, therefore we can write the two equations as
4s - 16 = 2s + 8
4s - 2s = 8 + 16
2s = 24
s = 12 yrs is the son
:
I'll you find the man's age using the 1st equation, check the solutions in the 2nd equation


find their present age